Grade 11Physical Chemistry

The Percentage of nitrogen in urea is about?A) 46. B)85C)18. D)28

Kashish
8 Years agoGrade 11

2 Answers

sankalp
8 Years ago
46%
 
urea=> NH2-CO-NH
total mass=60
mass of nitrogen = 14, 2 molecules hence m=28
N%= mass of nitrogen X 100
         mass of urea
  = 28  X 100
     60
  N%= 46.6%
Kushagra Madhukar
6 Years ago
Dear student,
Please find the attached solution to your problem. 
Molecular formula of urea is NH2-CO-NH
Hence, Molar mass = 60g
mass of nitrogen in 1 mole of Urea = 14*2 = 28g
Hence, percentage of Nitrogen = mass of nitrogen/mass of urea X 100
= 28/60 x 100
 = 46.6%
 
Hope it helps
Thanks and regards,
Kushagra